# Exactly solvable dissipative dynamics and one-form strong-to-weak spontaneous symmetry breaking in interacting two-dimensional spin systems

### v1

@@ IV.3 Gauge invariance
Because of the color rule, the edge operators (29) all commute with each other and with the Lindbladian. Moreover, they satisfy $u^{2}_{ij}=\tilde{u}^{2}_{ij}=v_{j}^{2}=1$. Hence, they can be viewed as classical $\mathbb{Z}_{2}$ gauge fields, $u_{ij},\tilde{u}_{ij},v_{i}\in\{-1,+1\}$, and we can replace the operators by these values in the Lindbladian. Since for each ±$\pm 1$ value of $u_{ij}$, $\tilde{u}_{ij}$, and $v_{i}$, Eq. (28) [or Eq. (35)] is quadratic in fermionic operators, the model is now exactly solvable. The Lindbladian block-diagonalizes into sectors where the gauge fields act as multiples of the identity:
=$\mathcal{L}=\bigoplus_{\{u_{ij},\tilde{u}_{ij},v_{j}\}}\mathcal{L}_{u_{ij},\tilde{u}_{ij},v_{j}}.$ (36)
@@
The total parity superoperator $\mathcal{D}=\prod_{j}D_{j}\widetilde{D}_{j}$ is a gauge transformation on every site, hence leaving all gauge fields invariant, and thus providing no gauge-field constraint. It however constrains the matter fields. Writing
@@
shows that the physical constraint $\mathcal{D}=+1$ can be seen as fixing the adjoint fermion parity to be the same as the “gauge parity” [35]: [^4] If $z(j)$ is odd, there are two additional Majoranas $b_{j}^{2k}$ and $\tilde{b}^{2k}$ that appear neither in the Lindbladian nor any gauge fields $u_{ij}$, $\tilde{u}_{ij}$, or $v_{j}$, but still enter the definition of $D_{j}$ and $\widetilde{D}_{j}$ and thus affect the constraint $\mathcal{D}=1$. Consequently, there is an additional factor in Eq. (41): $\prod^{\prime}_{j}(\mathrm{i}v^{2k}_{j})$, where $v^{2k}_{j}=\mathrm{i}b^{2k}_{j}\tilde{b}^{2k}_{j}$ (with $(v_{j}^{2k})^{2}=1$) and the product $\prod_{j}^{\prime}$ is over sites with odd $z(j)$. By the handshaking lemma, the number $2M$ of such sites is always even and hence $\prod^{\prime}_{j}(\mathrm{i}v^{2k}_{j})=(-1)^{M}\prod^{\prime}_{j}v^{2k}_{j}$. Since $v^{2k}_{j}$ does not enter the Lindbladian but commutes with it and with all gauge fields, we can simply choose an arbitrary gauge $v_{j}^{2k}=\pm 1$ such that $\prod^{\prime}_{j}v^{2k}_{j}=(-1)^{M}$ without changing the physics; Eq. (41) then holds for odd $z(j)$ as well.
@@
Since $D_{j}$ and $\widetilde{D}_{j}$ are gauge transformations, the physical subspace [defined by Eq. (26)] has gauge-invariant states. By Eqs. (37) and (38), the Lindbladian is also gauge invariant, as is its spectrum (because it is preserved by similarity transformations, such as gauge transformations). However, the eigenstates $|\varphi\rangle$ obtained from Eq. (28) [or (35)] are not, since we must choose some gauge to compute them (we fix this gauge in Sec. IV.4). To obtain physical eigenstates, we project $|\varphi\rangle$ into the physical subspace using the projector
@@
The physical state $\mathcal{Q}|\varphi\rangle$ corresponds to an equal-weight superposition of eigenstates for all gauge-equivalent configurations [^5] As there are only $N-1$ independent gauge transformations, despite its appearance $\mathcal{Q}$ does not perform $N$ independent projections, since Eq. (41) is already implemented (i.e., $\mathcal{Q}$ and $\mathcal{D}$ are not algebraically independent). $\{u_{ij},\tilde{u}_{ij},v_{j}\}$; these all have the same eigenvalue (which follows from the gauge invariance of the Lindbladian, $\mathcal{L}\mathcal{Q}|\varphi\rangle=\mathcal{Q}\mathcal{L}|\varphi\rangle$). $\mathcal{Q}|\varphi\rangle$ is the vectorization of $\mathcal{Q}(\varphi)=Q\varphi Q$, where $Q=\prod_{j}(\mathbbm{1}+D_{j})/2$; since $D_{j}Q=QD_{j}=Q$, $\mathcal{Q}(\varphi)$ satisfies Eq. (26) by construction.
A set of gauge-invariant observables that label physical states is furnished by string operators, but not all string operators: only the closed loops (flux operators) $W_{\mathcal{C}}$, $\widetilde{W}_{\mathcal{C}}$ and $\mathcal{W}_{\mathcal{P}}^{\mathrm{ch}}$ are gauge invariant. As mentioned in Sec. II.6 and shown in App. B, the elementary fluxes are the fluxes $W_{p}$ through the plaquettes in one layer, the fluxes $W_{\ell}$ through noncontractible loops in the same layer, and the fluxes $\mathcal{W}_{q}$ through interlayer plaquettes; their number totals $12E-N+$ 1. The fluxes $\widetilde{W}_{\mathcal{C}}$ in the second layer can be obtained from the preceding ones, $W_{\mathcal{C}}$ and $\mathcal{W}_{\mathcal{P}}^{\mathrm{ch}}$, since the flux through a closed cell is unity. This reasoning provides an alternative block-decomposition of $\mathcal{L}$ in terms of gauge-invariant fluxes,
@@
with each block labeled by a strong- and a weak-flux configuration $\bm{w}$ and $\bm{\mathfrak{w}}$, respectively. Since for each vertex $i$ there are $z(i)$ gamma matrices, by the handshaking lemma there are $\sum_{i}z(i)=2E$ gamma matrices in total. Subtracting the number $12E-N+$ 1 of (strong and weak) fluxes leaves $N-1$ binary degrees of freedom for the $c$ (and $\tilde{c}$) fermions [cf. the normalization $\mathcal{Z}$ of a sector with fixed fluxes in Eq. (13)], in agreement with there being one adjoint fermion mode per site and fixed total adjoint fermionic parity [Eq. (41)].
@@ IV.4 Gauge fixing
To perform actual calculations, we must fix a gauge. As discussed in Sec. III.1, there are exponentially many steady states, one for each strong-flux configuration $\bm{w}$ (i.e., vortex pattern) and $\bm{\mathfrak{w}}=\bm{1}$ (weak-symmetry constraint). The strong-symmetry constraint $W_{p}=\widetilde{W}_{p}$ is then automatically enforced as discussed in Sec. II.6 and App. B. To satisfy $\mathfrak{w}_{q}=1$ for all $q$ [cf. Eq. (31)] we make the gauge choice
@@
which is consistent with Eq. (34). [^6] This is a different gauge choice from that of Ref. 35, namely, $u_{ij}=-\tilde{u}_{ij}$ and staggered $v_{i}$ (note there is a minus sign difference between our definition of $\tilde{u}_{ij}$ and that of Ref. 35). While the two choices are equivalent for the square lattice, the gauge in Ref. 35 can only be generalized to bipartite graphs, whereas ours works for any graph. Moreover, this gauge choice allows us to completely gauge away the edge variables $v_{i}$, see below. For any $\bm{w}$, $\mathcal{L}|0\rangle_{\bm{w},\bm{\mathfrak{w}}=\bm{1}}=0$ and, hence, as expected from the discussion at the end of the previous subsection, we identify the adjoint-fermionic vacuum as the vectorization of the steady state $\rho_{\bm{w}}$. The excitations inside a steady-state sector are then obtained from the other $12^{N-1}-$ 1 eigenstates of $\mathcal{L}$ with $(-1)^{\mathcal{N}_{f}}=1$ [to maintain Eq. (41)].
@@
For a particular gauge choice $\{u_{ij}\}$ compatible with the background flux configuration $\bm{w}$ and set of gauge flips $\mathfrak{U}$ and $\mathfrak{V}$, the Lindbladian reads
@@
with the adjoint fermionic parity fixed to $(-1)^{\mathcal{N}_{f}}=(-1)^{|\mathfrak{U}|+|\mathfrak{V}|}$. Here, $\overline{\mathfrak{U}}$ ($\overline{\mathfrak{V}}$) denotes the complement of the set $\mathfrak{U}$ ($\mathfrak{V}$) and $|\mathfrak{U}|$ ($|\mathfrak{V}|$) the cardinality of $\mathfrak{U}$ ($\mathfrak{V}$).
From Eq. (46), we immediately see that excitations come in two types: if, for a given symmetry sector, there exists no gauge choice without intralayer flips ($\mathfrak{U}\neq\emptyset$), then there are superconducting terms in the Lindbladian and it conserves only adjoint-fermion-parity, $[\mathcal{L},(-1)^{\mathcal{N}_{f}}]=0$. On the other hand, if a gauge choice exists such that there are only interlayer flips or no gauge flips at all ($\mathfrak{U}=\emptyset$), then the superconducting terms vanish, and the dynamics conserves the adjoint-fermion number, $[\mathcal{L},\mathcal{N}_{f}]=0$. The projection into these symmetry sectors has more symmetry than the full Lindbladian.
Importantly, we can always make the gauge choice $\mathfrak{V}=\emptyset$ and perform all necessary gauge flips as intralayer gauge flips: for each $\mathfrak{w}_{q}=\mathfrak{w}_{\langle ij\rangle}=-1$, we can simply flip the $\tilde{u}_{ij}$ on the corresponding edge while keeping all $v_{j}=-1$ fixed (i.e., there is a one-to-one correspondence between $\tilde{u}_{ij}$ and $\mathfrak{w}_{\langle ij\rangle}$). The gauge choice $\mathfrak{V}=\emptyset$ would simplify the Lindbladian by dropping the last term in Eq. (46) and allow for a unified treatment of all sectors. However, as a consequence of their enhanced symmetry, sectors that are gauge equivalent to $\mathfrak{U}=\emptyset$ (i.e., where $W_{\mathcal{C}}=\widetilde{W}_{\mathcal{C}}$ for all $\mathcal{C}$) have a different phenomenology than those that are not. For this reason, we shall treat the two cases separately in Secs. V and VI, respectively. Before that, in Sec. IV.5, we identify the gamma-matrix strings in the original representation that correspond to the fractionalized excitations discussed above.
@@ IV.5 Steady states and excitations from strings
A generic state does not have definite quantum numbers under all conserved fluxes. To determine the relaxation time of such a state, we would need to find all the symmetry sectors with which it overlaps, and then compute the smallest spectral gap therein. This is an intractable task given the exponential number of sectors (for small system sizes, one can devise optimization schemes for this task [36] but, even then, the accessible system sizes are very limited).
An alternative is to compute the relaxation timescales of basis elements of the space of operators: any state can be written as a linear combination of terms of the form
@@
Crucially, any state of the form (47) is fully supported inside a single symmetry sector: $\rho_{0}$ has fixed and known quantum numbers, namely, $w_{p}=w_{\ell}=\mathfrak{w}_{q}=1$ for all $p$, $\ell$, and $q$; and all strings $\Sigma_{L}$ and $\Sigma_{R}$ either commute or anticommute with all conserved fluxes, hence either preserve or flip these quantum numbers. More precisely, if we denote
@@
with $s_{p},s_{\ell},\mathfrak{s}_{q}^{L},\mathfrak{s}_{q}^{R}\in\{-1,1\}$, then $\sigma$ belongs to the sector labeled by the quantum numbers $w_{p}=s_{p}$, $w_{\ell}=s_{\ell}$, and $\mathfrak{w}_{q}=\mathfrak{s}_{q}^{L}\mathfrak{s}_{q}^{R}$. We do not need to specify the commutation relation of $\Sigma_{R}$ with the strong fluxes, since $\widetilde{w}_{p}$ and $\widetilde{w}_{\ell}$ are completely fixed by $w_{p}$, $w_{\ell}$ and $\mathfrak{w}_{q}$. We then choose a gauge configuration $u_{ij}$ compatible with the strong fluxes and fix the inter- and intralayer gauge flips $\mathfrak{V}$ and $\mathfrak{U}$ (recall there is gauge freedom in this choice) in accordance with the weak fluxes.
@@
Our strategy is to choose strings $\Sigma_{L}$ and $\Sigma_{R}$ such that we can construct representative states with each type of excitation described in Sec. IV.4: vortices, intra- and interlayer gauge excitations, and Majorana fermions. To this end, it suffices to consider the strings $W^{\mu\nu}_{\mathcal{P}}$, $W_{\mathcal{P}}$, $W_{\mathcal{P}}^{\mathrm{ch}}$, and $W_{\mathcal{C}}$ defined in Eqs. (6)–(9), as described in the following and summarized in Fig. 4 and Table 1.
@@
    Open strings $W_{\mathcal{P}}^{\mu\nu}$ with $\mu,\nu\in\{1,\dots,2k\}$. $W_{\mathcal{P}}^{\mu\nu}$ is formed by a body $\prod_{\langle ij\rangle\in\mathcal{P}}K_{ij}$ and two gamma matrices at the endpoints. In the fractionalized picture, this maps to a string of gauge fields, $\prod_{\langle ij\rangle\in\mathcal{P}}(-\mathrm{i}u_{ij})$, with a Majorana fermion $b_{j}^{\mu}$ at the endpoint $j$ of $\mathcal{P}$ (and similarly for the other endpoint), see Fig. 4(a). We have $\mathfrak{s}_{q}^{L,R}=-1$ on the edge that emanates from $j$ with color $\mu$ and $s_{p}=-1$ on the two plaquettes adjacent to it, which in the fractionalized representation reads as the Majorana $b_{j}^{\mu}$ flipping the gauge field $u_{ij}$ on the edge with color $\mu$. If the action of $W_{\mathcal{P}}^{\mu\nu}$ is adjoint, the $\mathfrak{s}_{q}^{L,R}=-1$ cancel each other out ($\mathfrak{w}_{q}=\mathfrak{s}_{q}^{L}\mathfrak{s}_{q}^{R}=+1$) and the string excites only vortices in the system. Hence, $\sigma=W^{\mu\nu}_{\mathcal{P}}\rho_{0}W^{\mu\nu}_{\mathcal{P}}=P_{\bm{w}}/\mathcal{Z}$ is a different steady state, where $w_{p}=-1$ on the plaquettes mentioned above and $w_{p}=+1$ otherwise; in the vectorized representation, $\sigma$ maps to the vacuum $|0\rangle_{\bm{w}\neq\bm{1},\bm{\mathfrak{w}}=\bm{1}}$.
@@
    When $W^{\mu\nu}_{\mathcal{P}}$ acts on the right, we do not induce vortices ($s_{p}=+1$ for all $p$), but the $\mathfrak{s}_{q}^{L,R}$ no longer cancel, leading instead to an intralayer gauge excitation on the same edge ($\tilde{u}_{ij}=-u_{ij}$). The mapped state $\sigma=\rho_{0}W^{\mu\nu}_{\mathcal{P}}$ is not a steady steady state but instead a decaying operator in a symmetry sector with $w_{p}\neq\widetilde{w}_{p}$; its vectorization is the vacuum $|0\rangle_{\bm{w}=\bm{1},\bm{\mathfrak{w}}\neq\bm{1}}$. On the other hand, the left action can be interpreted as the composition of the right and adjoint actions. In that case, $\sigma$ has both vortices and intralayer gauge excitations.
@@
    Open chiral strings $W_{\mathcal{P}}^{\mathrm{ch}}$ ($\mu,\nu=2k+1$). Similarly to before, these strings have a “$\mathbb{Z}_{2}$ gauge body”, but carry chiral Majoranas $b_{j}^{\mathrm{ch}}$ at their endpoints, see Fig. 4(b). $W_{\mathcal{P}}^{\mathrm{ch}}$ commutes with all $W_{\mathcal{C}}$, thus $s_{p}=s_{\ell}=1$ for all $p$ and $\ell$, and chiral strings do not excite vortices. On the other hand, $\mathfrak{s}_{q}^{L,R}=-1$ for all edges $q$ that emanate from the endpoints of the string and $\mathfrak{s}_{q}^{L,R}=1$ for all other edges. In the fractionalized picture, it corresponds to the chiral Majorana $b_{j}^{\mathrm{ch}}$ flipping the gauge field $v_{j}$; this is gauge-equivalent to flipping all the gauge fields $u_{ij}$ emanating from site $j$ instead (as can also be seen by direct inspection of the commutation relations of $u_{ij}$, $\tilde{u}_{ij}$, $v_{j}$, and $b^{\mathrm{ch}}_{j}$), and we conclude that open chiral strings effect interlayer gauge flips at their endpoints. For the adjoint action, the interlayer gauge flips induced in each layer cancel and no excitations are induced. We thus see that $\sigma=W_{\mathcal{P}}^{\mathrm{ch}}\rho_{0}=\rho_{0}W_{\mathcal{P}}^{\mathrm{ch}}$ is a transient operator in a symmetry sector with $\mathfrak{w}_{q}=-1$ only and maps to a fermionic vacuum $|0\rangle_{\bm{w}=\bm{1},\bm{\mathfrak{w}}\neq\bm{1}}$ with $\bm{\mathfrak{w}}$ as described above. Hence, starting from the vacuum in the reference sector, we can generate all other possible vacua by acting with $W^{\mu\nu}_{\mathcal{P}}$ and $W_{\mathcal{P}}^{\mathrm{ch}}$.

### v2

@@ IV.3 Gauge invariance
Because of the color rule, the edge operators (29) all commute with each other and with the Lindbladian. Moreover, they satisfy $u^{2}_{ij}=\tilde{u}^{2}_{ij}=v_{j}^{2}=1$. Hence, they can be viewed as classical $\mathbb{Z}_{2}$ gauge fields, $u_{ij},\tilde{u}_{ij},v_{i}\in\{-1,+1\}$, and we can replace the operators by these values in the Lindbladian. Since for each $\pm 1$ value of $u_{ij}$, $\tilde{u}_{ij}$, and $v_{i}$, Eq. (28) [or Eq. (35)] is quadratic in fermionic operators, the model is now exactly solvable. The Lindbladian block-diagonalizes into sectors where the gauge fields act as multiples of the identity:
$\mathcal{L}=\bigoplus_{\{u_{ij},\tilde{u}_{ij},v_{j}\}}\mathcal{L}_{u_{ij},\tilde{u}_{ij},v_{j}}.$ (36)
@@
The total parity superoperator $\mathcal{D}=\prod_{j}D_{j}\widetilde{D}_{j}$ is a gauge transformation on every site, hence leaving all gauge fields invariant, and thus providing no gauge-field constraint. However, it constrains the matter fields. Writing
@@
shows that the physical constraint $\mathcal{D}=+1$ can be seen as fixing the adjoint fermion parity to be the same as the “gauge parity” [35]: [^4] If $z(j)$ is odd, there are two additional Majoranas $b_{j}^{2k}$ and $\tilde{b}_{j}^{2k}$ that appear neither in the Lindbladian nor any gauge fields $u_{ij}$, $\tilde{u}_{ij}$, or $v_{j}$, but still enter the definition of $D_{j}$ and $\widetilde{D}_{j}$ and thus affect the constraint $\mathcal{D}=1$. Consequently, there is an additional factor in Eq. (41): $\prod^{\prime}_{j}(\mathrm{i}v^{2k}_{j})$, where $v^{2k}_{j}=\mathrm{i}b^{2k}_{j}\tilde{b}^{2k}_{j}$ (with $(v_{j}^{2k})^{2}=1$) and the product $\prod_{j}^{\prime}$ is over sites with odd $z(j)$. By the handshaking lemma, the number $2M$ of such sites is always even and hence $\prod^{\prime}_{j}(\mathrm{i}v^{2k}_{j})=(-1)^{M}\prod^{\prime}_{j}v^{2k}_{j}$. Since $v^{2k}_{j}$ does not enter the Lindbladian but commutes with it and with all gauge fields, we can simply choose an arbitrary gauge $v_{j}^{2k}=\pm 1$ such that $\prod^{\prime}_{j}v^{2k}_{j}=(-1)^{M}$ without changing the physics; Eq. (41) then holds for odd $z(j)$ as well.
@@
Since $D_{j}$ and $\widetilde{D}_{j}$ are gauge transformations, the physical subspace [defined by Eq. (26)] has gauge-invariant states. By Eqs. (37) and (38), the Lindbladian is also gauge invariant, as is its spectrum (because it is preserved by similarity transformations, such as gauge transformations). However, the eigenstates $|\varphi\rangle$ obtained from Eq. (28) [or (35)] are not, since we must choose some gauge to compute them (we fix this gauge in Sec. IV.4). To obtain physical eigenstates, we project $|\varphi\rangle$ into the physical subspace using the projector [31, 35]
@@
The physical state $\mathcal{Q}|\varphi\rangle$ corresponds to an equal-weight superposition of eigenstates for all gauge-equivalent configurations [^5] As there are only N1N-1 independent gauge transformations, despite its appearance $\mathcal{Q}$ does not perform $N$ independent projections, since Eq. (41) is already implemented (i.e., $\mathcal{Q}$ and $\mathcal{D}$ are not algebraically independent). $\{u_{ij},\tilde{u}_{ij},v_{j}\}$; these all have the same eigenvalue (which follows from the gauge invariance of the Lindbladian, $\mathcal{L}\mathcal{Q}|\varphi\rangle=\mathcal{Q}\mathcal{L}|\varphi\rangle$). $\mathcal{Q}|\varphi\rangle$ is the vectorization of $\mathcal{Q}(\varphi)=Q\varphi Q$, where $Q=\prod_{j}(\mathbbm{1}+D_{j})/2$; since $D_{j}Q=QD_{j}=Q$, $\mathcal{Q}(\varphi)$ satisfies Eq. (26) by construction.
A set of gauge-invariant observables that label physical states is furnished by string operators, but not all string operators: only the closed loops (flux operators) $W_{\mathcal{C}}$, $\widetilde{W}_{\mathcal{C}}$, and $\mathcal{W}_{\mathcal{P}}^{\mathrm{ch}}$ are gauge invariant. As mentioned in Sec. II.6 and shown in Appendix B, the elementary fluxes are the fluxes $W_{p}$ through the plaquettes in one layer, the fluxes $W_{\ell}$ through noncontractible loops in the same layer, and the fluxes $\mathcal{W}_{q}$ through interlayer plaquettes; their number totals 2EN+12E-N+1. The fluxes $\widetilde{W}_{\mathcal{C}}$ in the second layer can be obtained from the preceding ones, $W_{\mathcal{C}}$ and $\mathcal{W}_{\mathcal{P}}^{\mathrm{ch}}$, since the flux through a closed cell is unity. This reasoning provides an alternative block-decomposition of $\mathcal{L}$ in terms of gauge-invariant fluxes,
@@
with each block labeled by a strong- and a weak-flux configuration $\bm{w}$ and $\bm{\mathfrak{w}}$, respectively. Since for each vertex $i$ there are $z(i)$ gamma matrices, by the handshaking lemma there are $\sum_{i}z(i)=2E$ gamma matrices in total. Subtracting the number 2EN+12E-N+1 of (strong and weak) fluxes leaves N1N-1 binary degrees of freedom for the $c$ (and $\tilde{c}$) fermions [cf. the normalization $\mathcal{Z}$ of a sector with fixed fluxes in Eq. (13)], in agreement with there being one adjoint fermion mode per site and fixed total adjoint fermionic parity [Eq. (41)].
@@ IV.4 Gauge fixing
To perform actual calculations, we must fix a gauge. As discussed in Sec. III.1, there are exponentially many steady states, one for each strong-flux configuration $\bm{w}$ (i.e., vortex pattern) and $\bm{\mathfrak{w}}=\bm{1}$ (weak-symmetry constraint). The strong-symmetry constraint $W_{p}=\widetilde{W}_{p}$ is then automatically enforced as discussed in Sec. II.6 and Appendix B. To satisfy $\mathfrak{w}_{q}=1$ for all $q$ [cf. Eq. (31)] we make the gauge choice
@@
which is consistent with Eq. (34). [^6] This is a different gauge choice from that of Ref. 35, namely, $u_{ij}=-\tilde{u}_{ij}$ and staggered $v_{i}$ (note there is a minus sign difference between our definition of $\tilde{u}_{ij}$ and that of Ref. 35). While the two choices are equivalent for the square lattice, the gauge in Ref. 35 can only be generalized to bipartite graphs, whereas ours works for any graph. Moreover, this gauge choice allows us to completely gauge away the edge variables $v_{i}$, see below. For any $\bm{w}$, $\mathcal{L}|0\rangle_{\bm{w},\bm{\mathfrak{w}}=\bm{1}}=0$ and, hence, as expected from the discussion at the end of the previous subsection, we identify the adjoint-fermionic vacuum as the vectorization of the steady state $\rho_{\bm{w}}$. The excitations inside a steady-state sector are then obtained from the other $2^{N-1}-1$ eigenstates of $\mathcal{L}$ with $(-1)^{\mathcal{N}_{f}}=1$ [to maintain Eq. (41)].
@@
For a particular gauge choice $\{u_{ij}\}$ compatible with the background flux configuration $\bm{w}$ and the set of gauge flips $\mathfrak{U}$ and $\mathfrak{V}$, the Lindbladian reads
@@
with the adjoint fermionic parity fixed to $(-1)^{\mathcal{N}_{f}}=(-1)^{|\mathfrak{U}|+|\mathfrak{V}|}$. Here, $\overline{\mathfrak{U}}$ ($\overline{\mathfrak{V}}$) denotes the complement of the set $\mathfrak{U}$ ($\mathfrak{V}$) and $|\mathfrak{U}|$ ($|\mathfrak{V}|$) denotes the cardinality of $\mathfrak{U}$ ($\mathfrak{V}$).
From Eq. (46), we immediately see that excitations come in two types: if, for a given symmetry sector, there exists no gauge choice without intralayer flips ($\mathfrak{U}\neq\emptyset$), then there are superconducting terms in the Lindbladian and it conserves only adjoint-fermion-parity, $\mathcal{L},(-1)^{\mathcal{N}_{f}}]=0$. On the other hand, if a gauge choice exists such that there are only interlayer flips or no gauge flips at all ($\mathfrak{U}=\emptyset$), then the superconducting terms vanish, and the dynamics conserves the adjoint-fermion number, $\mathcal{L},\mathcal{N}_{f}]=0$. The projection into these symmetry sectors has more symmetry than the full Lindbladian.
Importantly, we can always make the gauge choice $\mathfrak{V}=\emptyset$ and perform all necessary gauge flips as intralayer gauge flips: for each $\mathfrak{w}_{q}=\mathfrak{w}_{\langle ij\rangle}=-1$, we can simply flip the $\tilde{u}_{ij}$ on the corresponding edge while keeping all $v_{j}=-1$ fixed (i.e., there is a one-to-one correspondence between $\tilde{u}_{ij}$ and $\mathfrak{w}_{\langle ij\rangle}$). The gauge choice $\mathfrak{V}=\emptyset$ would simplify the Lindbladian by dropping the last term in Eq. (46) and allow for a unified treatment of all sectors. However, as a consequence of their enhanced symmetry, sectors that are gauge equivalent to $\mathfrak{U}=\emptyset$ (i.e., where $W_{\mathcal{C}}=\widetilde{W}_{\mathcal{C}}$ for all $\mathcal{C}$) have a different phenomenology from those that are not. For this reason, we shall treat the two cases separately in Secs. V and VI, respectively. Before that, in Sec. IV.5, we identify the gamma-matrix strings in the original representation that correspond to the fractionalized excitations discussed above.
@@ IV.5 Steady states and excitations from strings
A generic state does not have definite quantum numbers under all conserved fluxes. To determine the relaxation time of such a state, we would need to find all the symmetry sectors with which it overlaps, and then compute the smallest spectral gap therein. This is an intractable task given the exponential number of sectors (for small system sizes, one can devise optimization schemes for this task [36], but even then, the accessible system sizes are very limited).
An alternative is to compute the relaxation timescales of basis elements of the space of operators: any operator can be written as a linear combination of terms of the form
@@
Crucially, any state of the form (47) is fully supported inside a single symmetry sector: $\rho_{0}$ has fixed and known quantum numbers, namely, $w_{p}=w_{\ell}=\mathfrak{w}_{q}=1$ for all $p$, $\ell$, and $q$; and all strings $\Sigma_{L}$ and $\Sigma_{R}$ either commute or anticommute with all conserved fluxes, hence they either preserve or flip these quantum numbers. More precisely, if we denote
@@
with $s_{p},s_{\ell},\mathfrak{s}_{q}^{L},\mathfrak{s}_{q}^{R}\in\{-1,1\}$, then $\sigma$ belongs to the sector labeled by the quantum numbers $w_{p}=s_{p}$, $w_{\ell}=s_{\ell}$, and $\mathfrak{w}_{q}=\mathfrak{s}_{q}^{L}\mathfrak{s}_{q}^{R}$. We do not need to specify the commutation relation of $\Sigma_{R}$ with the strong fluxes, since $\widetilde{w}_{p}$ and $\widetilde{w}_{\ell}$ are completely fixed by $_{p}, ww_{\ell}$ and $\mathfrak{w}_{q}$. We then choose a gauge configuration $u_{ij}$ compatible with the strong fluxes and fix the inter- and intralayer gauge flips $\mathfrak{V}$ and $\mathfrak{U}$ (recall that there is gauge freedom in this choice) in accordance with the weak fluxes.
@@
Our strategy is to choose strings $\Sigma_{L}$ and $\Sigma_{R}$ such that we can construct representative states with each type of excitation described in Sec. IV.4: vortices, intra- and interlayer gauge excitations, and Majorana fermions. To this end, it suffices to consider the strings $W^{\mu\nu}_{\mathcal{P}}$, $W_{\mathcal{P}}$, $W_{\mathcal{P}}^{\mathrm{ch}}$, and $W_{\mathcal{C}}$ defined in Eqs. (6)–(9), as described in the following and summarized in Table 1 and Fig. 4.
@@
    Open strings $W_{\mathcal{P}}^{\mu\nu}$ with $\mu,\nu\in\{1,\dots,2k\}$. $W_{\mathcal{P}}^{\mu\nu}$ is formed by a body $\prod_{\langle ij\rangle\in\mathcal{P}}K_{ij}$ and two gamma matrices at the endpoints. In the fractionalized picture, this maps to a string of gauge fields, $\prod_{\langle ij\rangle\in\mathcal{P}}(-\mathrm{i}u_{ij})$, with a Majorana fermion $b_{j}^{\mu}$ at the endpoint $j$ of $\mathcal{P}$ (and similarly for the other endpoint); see Fig. 4(a). We have $\mathfrak{s}_{q}^{L,R}=-1$ on the edge that emanates from $j$ with color $\mu$, and $s_{p}=-1$ on the two plaquettes adjacent to it, which in the fractionalized representation reads as the Majorana $b_{j}^{\mu}$ flipping the gauge field $u_{ij}$ on the edge with color $\mu$. If the action of $W_{\mathcal{P}}^{\mu\nu}$ is adjoint, the $\mathfrak{s}_{q}^{L,R}=-1$ cancel each other out ($\mathfrak{w}_{q}=\mathfrak{s}_{q}^{L}\mathfrak{s}_{q}^{R}=+1$) and the string excites only vortices in the system. Hence, $\sigma=W^{\mu\nu}_{\mathcal{P}}\rho_{0}W^{\mu\nu}_{\mathcal{P}}=P_{\bm{w}}/\mathcal{Z}$ is a different steady state, where $w_{p}=-1$ on the plaquettes mentioned above and $w_{p}=+1$ otherwise; in the vectorized representation, $\sigma$ maps to the vacuum $|0\rangle_{\bm{w}\neq\bm{1},\bm{\mathfrak{w}}=\bm{1}}$.
@@
    When $W^{\mu\nu}_{\mathcal{P}}$ acts on the right, we do not induce vortices ($s_{p}=+1$ for all $p$), but the $\mathfrak{s}_{q}^{L,R}$ no longer cancel, leading instead to an intralayer gauge excitation on the same edge ($\tilde{u}_{ij}=-u_{ij}$). The mapped state $\sigma=\rho_{0}W^{\mu\nu}_{\mathcal{P}}$ is not a steady state but instead a decaying operator in a symmetry sector with $w_{p}\neq\widetilde{w}_{p}$; its vectorization is the vacuum $|0\rangle_{\bm{w}=\bm{1},\bm{\mathfrak{w}}\neq\bm{1}}$. On the other hand, the left action can be interpreted as the composition of the right and adjoint actions. In that case, $\sigma$ has both vortices and intralayer gauge excitations.
@@
    Open chiral strings $W_{\mathcal{P}}^{\mathrm{ch}}$ ($\mu,\nu=2k+1$). Similarly to before, these strings have a $\mathbb{Z}_{2}$ gauge body”, but they carry chiral Majoranas $b_{j}^{\mathrm{ch}}$ at their endpoints; see Fig. 4(b). $W_{\mathcal{P}}^{\mathrm{ch}}$ commutes with all $W_{\mathcal{C}}$, thus $s_{p}=s_{\ell}=1$ for all $p$ and $\ell$, and chiral strings do not excite vortices. On the other hand, $\mathfrak{s}_{q}^{L,R}=-1$ for all edges $q$ that emanate from the endpoints of the string and $\mathfrak{s}_{q}^{L,R}=1$ for all other edges. In the fractionalized picture, it corresponds to the chiral Majorana $b_{j}^{\mathrm{ch}}$ flipping the gauge field $v_{j}$; this is gauge-equivalent to flipping all the gauge fields $u_{ij}$ emanating from site $j$ instead (as can also be seen by direct inspection of the commutation relations of $u_{ij}$, $\tilde{u}_{ij}$, $v_{j}$, and $b^{\mathrm{ch}}_{j}$), and we conclude that open chiral strings effect interlayer gauge flips at their endpoints. For the adjoint action, the interlayer gauge flips induced in each layer cancel, and no excitations are induced. We thus see that $\sigma=W_{\mathcal{P}}^{\mathrm{ch}}\rho_{0}=\rho_{0}W_{\mathcal{P}}^{\mathrm{ch}}$ is a transient operator in a symmetry sector with $\mathfrak{w}_{q}=-1$ only and maps to a fermionic vacuum $|0\rangle_{\bm{w}=\bm{1},\bm{\mathfrak{w}}\neq\bm{1}}$ with $\bm{\mathfrak{w}}$ as described above. Hence, starting from the vacuum in the reference sector, we can generate all other possible vacua by acting with $W^{\mu\nu}_{\mathcal{P}}$ and $W_{\mathcal{P}}^{\mathrm{ch}}$.